# Elliptic curves and Fourier coefficients of meromorphic modular forms

### v1

@@ 4. The CM case and magnetic modular forms
We will now prove 2.1 when $p\geq 5$, or equivalently, $p\nmid 6$. As we will work with general number fields, we first introduce some notations.
@@
Let $L$ be a number field. For a prime $\mathfrak{p}$ of $L$, let $\mathbb{F}_{\mathfrak{p}}$ denote its residue field and let $\mathrm{N}(\mathfrak{p}):=|\mathbb{F}_{\mathfrak{p}}|$ denote its norm. For an elliptic curve $C/L$ and a good prime $\mathfrak{p}$ of $C$, define
@@
Now, for $k\in\{4,6,8,10,14\}$ and an elliptic curve $C/L$, write
$F_{k,C}\;:=\;\frac{E_{k}}{j-j(C)}.$
For the convenience of the discussion in this section, we also write
$F_{k,c}\;:=\;\frac{E_{k}}{j-c},$
for $k\in\{4,6,8,10,14\}$ and $c\in\overline{\mathbb{Q}}$ or $\overline{\mathbb{Z}_{p}}$. The main theorem that we will prove in this section is the following.
@@
Let $k\in\{4,6,8,10,14\}$, $L$ be a number field, and $C/L$ be an elliptic curve. Let $\mathfrak{p}$ be a good prime of $C$ with $\mathfrak{p}\nmid 6$ and $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,a
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{k,C})\;\equiv\;a_{\mathfrak{p}}(C)^{k-2}a_{n}(F_{k,C})\pmod{\mathfrak{p}}$.
As mentioned in Remark 2.2, the proof depends on the hypergeometric interpretation of the coefficients in the case when $k=4$. In particular, we will show the following result.
@@
$a_{p^{l}}(F_{4,c})$ $\;\equiv\;\big(c(c-1728)\big)^{\frac{p^{l}-1}{2}}\cdot\sum_{m=0}^{p^{l}-1}\frac{(6m)!}{(m!)^{3}(3m)!}{\hskip 1.0pt}c^{-m}\pmod{p}$.
We start by first recalling the definitions of the Hecke operators (on $q$-series). For $k\in\mathbb{Z}$, a rational prime $p$, and a $q$-series $f=\sum_{n\gg-\infty}a_{n}(f)q^{n}$, define
More generally, for a $q$-series $f=\sum_{n\gg-\infty}a_{n}(f)q^{n}$, $k\in\mathbb{Z}$, and $m\in\mathbb{Z}^{+}$, define $f|T_{m,k}$ by setting
$a_{n}(f|T_{m,k})\;:=\;\sum_{r>0,r|(m,n)}r^{k-1}a_{mn/r^{2}}(f).$
It is easy to check that for $m,n\in\mathbb{Z}^{+}$ and $k\in\mathbb{Z}$,
$T_{m,k}{\hskip 1.0pt}T_{n,k}\;=\;\sum_{r>0,r|(m,n)}r^{k-1}T_{mn/r^{2},k}.$
5.1. A first reinterpretation
For each $k,n$, there exists a polynomial $P_{k,n}\in\mathbb{Z}[X]$ of degree n1n-1 such that
$\frac{E_{k}}{j-X}\;=\;\sum_{n=1}^{\infty}P_{k,n}(X)q^{n}.$
Then,
(8) $a_{n}(F_{k,c})\;=\;P_{k,n}(c).$
@@ 5.1. A first reinterpretation
For $k\in\{4,6,8,10,14\}$,
(9) $\frac{n^{k-1}(E_{14-k}/\Delta)|T_{n,2-k}}{E_{14-k}/\Delta}\;=\;P_{k,n}(j),$
where $E_{0}:=1$.
@@
By [AKN97, Theorem 3] or [DJ08a, Theorem 2], it is known that
$G_{k}(z,\tau)\;=\;\frac{E_{k}(z)\cdot(E_{14-k}/\Delta)(\tau)}{j(z)-j(\tau)}\;=\;\sum_{n=1}^{\infty}g_{2-k,n}(\tau)e^{2\pi inz}$
with $g_{2-k,n}=n^{k-1}(E_{14-k}/\Delta)|T_{n,2-k}$. The lemma then follows from dividing both sides of the expansion by $(E_{14-k}/\Delta)(\tau)$. ∎
@@
Let $k\in 2\mathbb{Z}^{+}$, $p$ be a rational prime, and $g\;=\;\sum_{m\gg-\infty}a_{m}(g)q^{m}$ with $a_{m}(g)\in\overline{\mathbb{Z}_{p}}$. For each $n\in\mathbb{Z}^{+}$, let
$g_{n}\;:=\;n^{k-1}g|T_{n,2-k}.$
Then,
$g_{np}\;\equiv\;g_{n}|V_{p}\pmod{p^{k-1}}$
as $q$-series.
Proof.
Since $T_{np,2-k}=T_{n,2-k}T_{p,2-k}-p^{1-k}T_{n/p,2-k}$, where $T_{n/p,2-k}:=0$ if $p\nmid n$, it follows that
$g_{np}\;=\;p^{k-1}g_{n}|T_{p,2-k}-p^{k-1}g_{n/p}\;=\;g_{n}|V_{p}+p^{k-1}g_{n}|U_{p}-p^{k-1}g_{n/p}\;\equiv\;g_{n}|V_{p}\pmod{p^{k-1}}$,
where $g_{n/p}:=0$ if $p\nmid n$. The result then follows. ∎
Corollary 5.5.
Let $k\in\{4,6,8,10,14\}$ and $p$ be a rational prime. Let $g_{2-k}=E_{14-k}/\Delta$ and let $g_{2-k,n}=n^{k-1}g_{2-k}|T_{n,2-k}$. Then,
$P_{k,np}(j)\;\equiv\;\frac{g_{2-k,n}^{p}}{g_{2-k}}\pmod{p}$
as $q$-series. In particular, for $l\in\mathbb{Z}^{+}$,
(10) $P_{k,p^{l}}(j)\;\equiv\;g_{2-k}^{p^{l}-1}\pmod{p}$
as $q$-series.
Proof.
Indeed,
$P_{k,np}(j)\;=\;\frac{g_{2-k,np}}{g_{2-k}}\;\equiv\;\frac{g_{2-k,n}|V_{p}}{g_{2-k}}\;\equiv\;\frac{g_{2-k,n}^{p}}{g_{2-k}}\pmod{p}$,
where we also use the fact that $g_{2-k,n}$ has integer coefficients. The second congruence now follows from induction. ∎
Let $n\in\mathbb{Z}^{+}$ and $p$ be a rational prime. Then, by Corollary 5.5
$\frac{P_{k,np}(j)}{P_{k,p}(j)P_{k,n}(j)}\;\equiv\;\frac{g_{2-k,n}^{p}/g_{2-k}}{(g_{2-k}^{p}/g_{2-k})\cdot(g_{2-k,n}/g_{2-k})}\;\equiv\;\bigg(\frac{g_{2-k,n}}{g_{2-k}}\bigg)^{p-1}\;\equiv\;P_{k,n}(j)^{p-1}\pmod{p}$
as $q$-series. Hence,
$P_{k,np}(j)\;\equiv\;P_{k,p}(j)\cdot P_{k,n}(j)^{p}\pmod{p}$
as polynomials in $j$. Plugging in $j=c\in\overline{\mathbb{Z}_{p}}$ and using that $a_{n}(F_{k,c})=P_{k,n}(c)$, we thus obtain that
$a_{np}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})a_{n}(F_{k,c})^{p}\pmod{p}$.
For $l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,c})\;\equiv\;a_{p}(F_{k,c}){\hskip 1.0pt}a_{np^{l-1}}(F_{k,c})^{p}\;\equiv\;\cdots\;\equiv\;a_{p}(F_{k,c})^{\frac{p^{l}-1}{p-1}}{\hskip 1.0pt}a_{n}(F_{k,c})^{p^{l}}\pmod{p}$.
Now, suppose that $L$ is a number field and $c\in L$. Then, for any prime $\mathfrak{p}$ of $L$ with $v_{\mathfrak{p}}(c)\geq 0$,
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})^{\frac{\mathrm{N}(\mathfrak{p})-1}{p-1}}a_{n}(F_{k,c})\pmod{\mathfrak{p}}$.
Since $a_{1}(F_{k,c})=1$, this also implies that
$a_{\mathrm{N}(\mathfrak{p})}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})^{\frac{\mathrm{N}(\mathfrak{p})-1}{p-1}}\pmod{\mathfrak{p}}$.
In summary, we obtain the following propositions.
Proposition 5.6.
Let $k\in\{4,6,8,10,14\}$, $p$ be a rational prime, and $c\in\overline{\mathbb{Z}_{p}}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})^{\frac{p^{l}-1}{p-1}}{\hskip 1.0pt}a_{n}(F_{k,c})^{p^{l}}\pmod{p}$.
Proposition 5.7.
Let $k\in\{4,6,8,10,14\}$, $L$ be a number field, $c\in L$, and $\mathfrak{p}$ be a prime of $L$ with $v_{\mathfrak{p}}(c)\geq 0$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})^{\frac{\mathrm{N}(\mathfrak{p})-1}{p-1}}a_{n}(F_{k,c})\;\equiv\;a_{\mathrm{N}(\mathfrak{p})}(F_{k,c}){\hskip 1.0pt}a_{n}(F_{k,c})\pmod{\mathfrak{p}}$.
From this, to prove Theorem 5.1, it suffices to show that
$a_{\mathrm{N}(\mathfrak{p})}(F_{k,C})\;\equiv\;a_{\mathfrak{p}}(C)^{k-2}\pmod{\mathfrak{p}}$.
5.2. Reducing to $k=4$
We first prove that all the cases when $k\in\{4,6,8,10,14\}$ in Theorem 5.1 follow from the single case when $k=4$.
Let $p$ be a rational prime and let $l\in\mathbb{Z}^{+}$. By Corollary 5.5,
$P_{k,p^{l}}(j)\;\equiv\;(E_{14-k}/\Delta)^{p^{l}-1}\pmod{p}$
as $q$-series. For $k\in\{6,8,10,14\}$, note that
[…]
Hence, for $k\in\{6,8,10,14\}$,
$P_{4,p^{l}}(j)^{\frac{k-2}{2}}\;\equiv\;\big((E_{10}/\Delta)^{\frac{k-2}{2}}\big)^{p^{l}-1}$ $\;=\;\big(j^{a}(j-1728)^{b}\big)^{p^{l}-1}(E_{14-k}/\Delta)^{p^{l}-1}$
$\;\equiv\;\big(j^{a}(j-1728)^{b}\big)^{p^{l}-1}P_{k,p^{l}}(j)\pmod{p}$,
as $q$-series, for some $a,b\in\mathbb{Z}_{\geq 0}$. This implies that
(11) $P_{4,p^{l}}(j)^{\frac{k-2}{2}}\;\equiv\;\big(j^{a}(j-1728)^{b}\big)^{p^{l}-1}P_{k,p^{l}}(j)\pmod{p}$
as polynomials in $j$. We thus obtain the following result.
Lemma 5.8.
Let $L$ be a number field and $C/L$ be an elliptic curve. Let $\mathfrak{p}$ be a prime of $L$ with $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. If $a_{\mathrm{N}(\mathfrak{p})}(F_{4,C})\equiv a_{\mathfrak{p}}(C)^{2}\pmod{\mathfrak{p}}$, then $a_{\mathrm{N}(\mathfrak{p})}(F_{k,C})\equiv a_{\mathfrak{p}}(C)^{k-2}\pmod{\mathfrak{p}}$ for all $k\in\{4,6,8,10,14\}$.
Proof.
Let $p$ be the rational prime lying below $\mathfrak{p}$ and write $\mathrm{N}(\mathfrak{p})=p^{l}$. For $k\in\{6,8,10,14\}$, by Equations (8) and (11), there exist $a,b\in\mathbb{Z}_{\geq 0}$ such that
(12) $a_{p^{l}}(F_{4,C})^{\frac{k-2}{2}}\;\equiv\;\big(j(C)^{a}(j(C)-1728)^{b}\big)^{p^{l}-1}a_{p^{l}}(F_{k,C})\;\equiv\;a_{p^{l}}(F_{k,C})\pmod{\mathfrak{p}}$.
The result then follows. ∎
@@ 5.3. Hypergeometric reinterpretation of $a_{p^{l}}(F_{4,c})$
Now, we specialize to $a_{p^{l}}(F_{4,c})$ and prove Theorem 5.2.
@@
Let $p\nmid 6$ be a rational prime and let $l\in\mathbb{Z}^{+}$. Recall the following identity by Fricke and Klein [Fri12, Chapter 5, Section 10]
@@
Now, write
@@
Here for a power series $R(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$, we write
$\mathrm{Trun}_{m}(R)(z)\;:=\;\sum_{n=0}^{m}a_{n}z^{n}\quad\text{ and }\quad\mathrm{Trun}_{m,z}\bigg(\sum_{n=0}^{\infty}a_{n}z^{n}\bigg)\;:=\;\sum_{n=0}^{m}a_{n}z^{n}.$
For $p\nmid 6$, it is easy to check that $\frac{(\frac{1}{2})_{r}(\frac{1}{6})_{r}(\frac{5}{6})_{r}}{(r!)^{3}}$ is divisible by $p$ for $\frac{p^{l}-1}{6}<r\leq p^{l}-1$. It follows that
$\mathrm{Trun}_{p^{l}-1,\alpha}\big((1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot B(\alpha)\big)\;=\;$ $\mathrm{Trun}_{p^{l}-1,\alpha}\big((1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{p^{l}-1}(B)(\alpha)\big)$
$\;\equiv\;$ $\mathrm{Trun}_{p^{l}-1,\alpha}\big((1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{\lfloor(p^{l}-1)/6\rfloor}(B)(\alpha)\big)$
$\;\equiv\;$ $(1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{\lfloor(p^{l}-1)/6\rfloor}(B)(\alpha)$
$\;\equiv\;$ $(1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{p^{l}-1}(B)(\alpha)\pmod{p}$.
@@
Let $L$ be a number field, $C/L$ be an elliptic curve, and $\mathfrak{p}$ be a prime of $L$ with $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Let $\mathbb{F}_{\mathfrak{p}}$ denote the residue field of $\mathfrak{p}$, let $p$ be the rational prime lying under $\mathfrak{p}$, and write $\mathrm{N}(\mathfrak{p})=p^{l}$. For $a\in L$ with $v_{\mathfrak{p}}(a)\geq 0$, define a quadratic character
@@
This congruence should follow from the hypergeometric machinery, though an explicit reference seems hard to find. As the proof is solely hypergeometric and unrelated to the main discussion of this paper, we will only write a sketch here.  [^8] The author plans to write a separate and self-contained note on the proof of this congruence in details.
@@
$C_{1}:y^{2}+xy\;=\;x^{3}-\frac{\lambda}{432}$
with $\lambda=\frac{1-\sqrt{1-1728/c}}{2}$ (so that $j(C_{1})=c$). It is known that for a prime $\mathfrak{P}$ of $L(\sqrt{1-1728/c})$ that is a good prime of $C_{1}$,

### v2

@@ 6.4. Negative discriminants
We will now prove Theorem 5.1. As we will work with general number fields, we first introduce some notations.
@@
Let $L$ be a number field and $\mathcal{O}_{L}$ be its ring of integers. For a prime $\mathfrak{p}$ of $L$, let $\mathbb{F}_{\mathfrak{p}}$ denote its residue field and let $\mathrm{N}(\mathfrak{p}):=|\mathbb{F}_{\mathfrak{p}}|$ denote its norm. For an elliptic curve $C/L$ and a good prime $\mathfrak{p}$ of $C$, define
@@
Now, for a holomorphic modular form $g\in\mathcal{O}_{L}[[q]]$ of weight $k$ and level $1$ and an elliptic curve $C/L$, write
$F_{g,C}\;:=\;\frac{g}{j-j(C)}.$
We also adopt the notation $g_{k}$ in Section 5, and for $k\in 2\mathbb{Z}_{\geq 2}$, write
$F_{k,C}\;:=\;F_{g_{k},C}\;=\;\frac{g_{k}}{j-j(C)}.$
The main theorem that we will prove in this section is the following.
@@
Let $k\in 2\mathbb{Z}_{\geq 2}$, $L$ be a number field, $g\in\mathcal{O}_{L}[[q]]$ be a holomorphic modular form of weight $k$ and level $1$, and $C/L$ be an elliptic curve. Let $\lambda$ be a relation for $S_{k}$ and let $\mathfrak{p}$ be a good prime of $C$ with $\mathfrak{p}\nmid 6$ and $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{g,C}|\lambda)\;\equiv\;a_{\mathfrak{p}}(C)^{k-2}a_{n}(F_{g,C}|\lambda)\pmod{\mathfrak{p}}$.
Remark 7.2.
By Lemma 5.5, it suffices to prove this for $g=g_{k}$, i.e., for $F_{k,C}$. Indeed, there exist some $\beta\in\mathcal{O}_{L}$ such that $F_{g,C}-\beta F_{k,C}\in S_{k}$. Hence, $F_{g,C}|\lambda=\beta\cdot F_{k,C}|\lambda$.
As mentioned in Remark 1.9, the proof depends on the hypergeometric interpretation of the coefficients in the case when $k=4$. In particular, we will show the following generalization of Theorem 1.6.
@@
$a_{p^{l}}\bigg(\frac{E_{4}}{j-c}\bigg)$ $\;\equiv\;\big(c(c-1728)\big)^{\frac{p^{l}-1}{2}}\cdot\sum_{m=0}^{p^{l}-1}\frac{(6m)!}{(m!)^{3}(3m)!}{\hskip 1.0pt}c^{-m}\pmod{p}$.
7.1. The two variable function
From now on, we fix $k\in 2\mathbb{Z}_{\geq 2}$, and write $k=k^{\prime}+12\delta$ for some $k^{\prime}\in\{0,4,6,8,10,14\}$ and $\delta\in\mathbb{Z}$. Recall that
$G_{k}(z,\tau)\;=\;\frac{g_{k}(z)\cdot g_{2-k}(\tau)}{j(z)-j(\tau)}.$
Let $q_{z}=e^{2\pi iz}$ and $q_{\tau}=e^{2\pi i\tau}$. Then, we have the following result by Duke–Jenkins, from which they deduced Zagier’s duality.
Theorem 7.4 ([DJ08a, Theorem 2]).
For any even integer $k\in\mathbb{Z}^{+}$,
$\sum_{n=\delta+1}^{\infty}g_{2-k,n}(\tau)q_{z}^{n}\;=\;G_{k}(z,\tau)\;=\;-\sum_{n=-\delta}^{\infty}g_{k,n}(z)q_{\tau}^{n}.$ (22)
Fix a rational prime $p$ and let $l\in\mathbb{Z}^{+}$. It is easy to check from the definitions that
$T_{p^{l},k}\;=\;\sum_{a=0}^{l}p^{a(k-1)}U_{p}^{l-a}V_{p}^{a}.$
Now, fix $n\in\mathbb{Z}$ with $n\geq-\delta$ and consider $g_{k,n}(z)|T_{p^{l},k}$. If $n\geq 0$, then we have
$g_{k,n}(z)\big|T_{p^{l},k}\;=\;\sum_{a=0}^{l}p^{a(k-1)}q_{z}^{-np^{2a-l}}+O(q^{\delta+1}),$
and it follows that
$g_{k,n}(z)\big|T_{p^{l},k}-\sum_{a=0}^{l}p^{a(k-1)}g_{k,np^{2a-l}}(z)\in S_{k},$ (23)
where we abuse the notations by setting that $q_{z}^{\alpha}:=0$ if $\alpha\notin\mathbb{Z}$, and that $g_{k,\alpha}:=0$ if $\alpha\notin\mathbb{Z}$, for convenience. If $\delta\leq n<0$, then we have
$g_{k,n}(z)\in S_{k}\quad\text{ and }\quad g_{k,n}(z)\big|T_{p^{l},k}\in S_{k}.$ (24)
Now, let $\lambda$ be a relation for $S_{k}$. In the following computations, the relation $\lambda$ and all the Hecke operators are viewed as acting on the variable $z$. We have
[…]
$\;\;\equiv\;-\sum_{n=0}^{\infty}\big(g_{k,np^{-s}}(z)\big|\lambda\big)q_{\tau}^{n}$
[…]
as a double Fourier series in terms of $z$ and $\tau$. Also,
$G_{k}(z,p^{l}\tau)\;=\;\frac{g_{k}(z)\cdot g_{2-k}(p^{l}\tau)}{j(z)-j(p^{l}\tau)}\;\equiv\;\frac{g_{k}(z)\cdot g_{2-k}(\tau)^{p^{l}}}{j(z)-j(\tau)^{p^{l}}}\pmod{p}$
as a double Fourier series in terms of $z$ and $\tau$. We thus obtain the following.
Proposition 7.5.
For a rational prime $p$, some $l\in\mathbb{Z}^{+}$, and a relation $\lambda$ for $S_{k}$, we have
$\frac{g_{k}(z)}{j(z)-j(\tau)}\bigg|\lambda U_{p^{l}}\;\equiv\;g_{2-k}(\tau)^{p^{l}-1}\cdot\frac{g_{k}(z)}{j(z)-j(\tau)^{p^{l}}}\bigg|\lambda\pmod{p}$
as a double Fourier series in terms of $z$ and $\tau$.
7.2. Reducing to the case of weight 44
We will now prove that all the cases follow from the single case when $k=4$ and $\lambda=(1,0,0,\ldots)$.
Write
$\frac{g_{k}}{j-X}\bigg|\lambda\;=\;\sum_{n=1}^{\infty}P_{k,n,\lambda}(X)q^{n}$
and in particular,
$\frac{E_{4}}{j-X}\;=\;\sum_{n=1}^{\infty}P_{4,n}(X)q^{n},$
for some polynomials $P_{k,n,\lambda},P_{4,n}\in\mathbb{Z}[X]$.
@@ 7.2. Reducing to the case of weight 44
For a rational prime $p$ and $l\in\mathbb{Z}^{+}$,
$P_{4,p^{l}}(j(\tau))\;\equiv\;(g_{-2}(\tau))^{p^{l}-1}\pmod{p}$
as $q_{\tau}$-series.
@@
This follows from specializing Proposition 7.5 at $k=4$ and $\lambda=(1,0,0,\ldots)$. ∎
Now, by analyzing the potential poles and zeros, it is easy to check that there always exist $a,b\in\mathbb{Z}_{\geq 0}$ such that
$\frac{g_{-2}(\tau)^{\frac{k-2}{2}}}{g_{2-k}(\tau)}\;=\;j(\tau)^{a}(j(\tau)-1728)^{b}.$
Then, Proposition 7.5 and Lemma 7.6 together imply that
$\bigg(\frac{g_{-2}(\tau)^{\frac{k-2}{2}}}{g_{2-k}(\tau)}\bigg)^{p^{l}-1}\cdot\frac{g_{k}(z)}{j(z)-j(\tau)}\bigg|\lambda U_{p^{l}}\;\equiv\;P_{4,p^{l}}(j(\tau))^{\frac{k-2}{2}}\cdot\frac{g_{k}(z)}{j(z)-j(\tau)^{p^{l}}}\bigg|\lambda\pmod{p}$
as a double Fourier series in terms of $z$ and $\tau$. In particular, taking the $n$-th coefficient of $q_{z}$ on both sides, one obtains that
$\big(j(\tau)^{a}(j(\tau)-1728)^{b}\big)^{p^{l}-1}P_{k,np^{l},\lambda}(j(\tau))\;\equiv\;P_{4,p^{l}}(j(\tau))^{\frac{k-2}{2}}P_{k,n,\lambda}(j(\tau)^{p^{l}})\pmod{p}$ (25)
as $q_{\tau}$-series, and hence also as polynomials in $j(\tau)$.
Now, let $L$ be a number field, $\mathfrak{p}$ be a prime of $L$, and $c\in L$ with $v_{\mathfrak{p}}(c)=0=v_{\mathfrak{p}}(c-1728)$. Let $p$ be a rational prime lying under $\mathfrak{p}$ and write $\mathrm{N}(\mathfrak{p})=p^{l}$. Then, plugging in $j(\tau)=c$ in (25) yields that
$P_{k,np^{l},\lambda}(c)\;\equiv\;P_{4,p^{l}}(c)^{\frac{k-2}{2}}P_{k,n,\lambda}(c)\pmod{\mathfrak{p}}$.
In particular, we obtain the following.
@@
Let $k\in 2\mathbb{Z}_{\geq 2}$, $L$ be a number field, and $C/L$ be an elliptic curve. Let $\lambda$ be a relation for $S_{k}$ and let $\mathfrak{p}$ be a prime with $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{k,C}|\lambda)\;\equiv\;a_{\mathrm{N}(\mathfrak{p})}(F_{4,C})^{\frac{k-2}{2}}{\hskip 1.0pt}a_{n}(F_{k,C}|\lambda)\pmod{\mathfrak{p}}$.
From the lemma above, it suffices to show that
$a_{\mathrm{N}(\mathfrak{p})}(F_{4,C})\;\equiv\;a_{\mathfrak{p}}(C)^{2}\pmod{\mathfrak{p}}$
in order to show Theorem 5.1.
@@ 7.3. Hypergeometric reinterpretation of $a_{p^{l}}(F_{4,C})$
Now, we specialize to $a_{p^{l}}(F_{4,C})$. Specifically, we will prove Theorem 7.3. For this part, it is more convenient to forget the dependence on the elliptic curves and simply write
$F_{4,c}\;:=\;\frac{E_{4}}{j-c},$
for $c\in\overline{\mathbb{Q}}$ or $\overline{\mathbb{Z}_{p}}$.
Recall that $a_{p^{l}}(F_{4,c})=P_{4,p^{l}}(c)$ and
$P_{4,p^{l}}(j)\;\equiv\;g_{-2}^{p^{l}-1}\;\equiv\;(E_{10}/\Delta)^{p^{l}-1}\pmod{p}$
as $q$-series. In order to prove Theorem 7.3, we will use the hypergeometric intepretation of $E_{4}$.
@@
Recall the following identity by Fricke and Klein [Fri12, Chapter 5, Section 10]
@@
Now, let $p\nmid 6$ be a rational prime and let $l\in\mathbb{Z}^{+}$. Write
@@
Here for a power series $R(t)=\sum_{n=0}^{\infty}a_{n}t^{n}$, we write
$\mathrm{Trun}_{m}(R)(t)\;:=\;\sum_{n=0}^{m}a_{n}t^{n}\quad\text{ and }\quad\mathrm{Trun}_{m,t}\bigg(\sum_{n=0}^{\infty}a_{n}t^{n}\bigg)\;:=\;\sum_{n=0}^{m}a_{n}t^{n}.$
For $p\nmid 6$, it is easy to check that $\frac{(\frac{1}{2})_{r}(\frac{1}{6})_{r}(\frac{5}{6})_{r}}{(r!)^{3}}$ is divisible by $p$ for $\frac{p^{l}-1}{6}<r\leq p^{l}-1$. From this, one can easily deduce that
$\mathrm{Trun}_{p^{l}-1,\alpha}\big((1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot B(\alpha)\big)\;\equiv\;(1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{p^{l}-1}(B)(\alpha)\pmod{p}$.
@@
Let $L$ be a number field, $C/L$ be an elliptic curve, and $\mathfrak{p}$ be a good prime of $C$ with $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Let $\mathbb{F}_{\mathfrak{p}}$ denote the residue field of $\mathfrak{p}$ and let $p$ be the rational prime lying under $\mathfrak{p}$. For $a\in L$ with $v_{\mathfrak{p}}(a)\geq 0$, define a quadratic character
@@
It is easy to check that if $\mathfrak{p}\nmid 2$, then
$\bigg(\frac{a}{\mathfrak{p}}\bigg)\;=\;a^{\frac{\mathrm{N}(\mathfrak{p})-1}{2}}\hskip 5.69046pt\text{ in }\mathbb{F}_{\mathfrak{p}}.$
@@
This congruence should follow from the hypergeometric machinery, though an explicit reference seems hard to find. Here we write a sketch of the proof and refer to the expository note [Zha25] for an explicit proof of this congruence.
@@
$C_{1}:\;y^{2}+xy\;=\;x^{3}-\frac{t}{432}$
with $t=\frac{1-\sqrt{1-1728/c}}{2}$ (so that $j(C_{1})=c$). It is known that for a prime $\mathfrak{P}$ of $L(\sqrt{1-1728/c})$ that is a good prime of $C_{1}$,